# LLMs Love Python: A Study of LLMs’ Bias for Programming Languages and Libraries

### v1

@@ 3.5.2. Rank Correlation Analysis
Table 3. Language Preferences, Benchmark Tasks (RQ1). Programming languages used by LLMs when solving problems from widely-used benchmarks. Column “Dataset” shows the name of the benchmark, as well as the programming languages for which the ground truth solutions are provided in each benchmark. For each LLM and dataset, the preferred languages (l) are given, along with the percentage (p) of dataset problems that have a solution in that language, and the total count of different languages used (when necessary). The most-used language in each case is in bold.
@@
4. Results
This section introduces the results. Note that it is expected that the percentages in the results do not add up to 100% because LLMs were prompted multiple times per task and sometimes responded with different languages or libraries. Additionally, the LLMs would sometimes provide a response without code; or would provide a response that contained code in multiple languages or that imported multiple libraries.
4.1. RQ1: Language Preference
@@ 4.1.1. Preferences for benchmark tasks
Table 3 shows the languages used by LLMs when solving problems from widely-used benchmarks, 200 problems were used per benchmark and each problem had three solutions generated. We observe a clear preference towards using Python for all LLMs and in all datasets.
For each dataset apart from AixBench, all LLMs gave solutions in Python for at least 93.5% of problems, the highest percentage of problems solved by a non-Python language was considerably lower at 19.5%. AixBench had a much more even distribution, with Python solutions given for only 53.5-78.3% of problems, with the second most-used language, Java, having solutions for 18.6-59.7% of problems. AixBench problems solved by the Claude models (Haiku3.5 and Sonnet3.5) are the only instances where Python was not the most-used language.
The “Real-world” datasets (AixBench and CoNaLa) had up to 12 different languages used in their solutions, significantly more than the other datasets - which had at most 5 different languages used, showing that models opted for more suitable languages in real-world situations. In general the range of programming languages that LLMs choose to use is limited, hundreds of programming languages get contributions on GitHub every year (TopProgrammingLanguages2022), but LLMs only choose to use 6-14 different ones.
We suspect this heavy preference for Python stems from LLMs being created with a focus on achieving state-of-the-art results on widely-used benchmarks, the vast majority of which are Python based, causing their training data to be saturated with Python code. When investigating what datasets to use for this study, it was clear that there was a lack of diversity in the datasets available and it was difficult to find suitable options with no link to Python. Whilst all datasets used have language-agnostic problem formations, some have been seeded from Python (Multi-HumanEval, MBXP, CoNaLa), or have Python solutions published (APPS, CodeContests). AixBench is the only dataset with no link to Python, which is likely why it is the closest to favouring another language.
Table 4. Language Preferences, Project Initialisation Tasks (RQ1). Programming languages chosen by LLMs for initial project code in scenarios where Python is not the most suitable option. For each LLM and project, each language (l) used by the LLM is given, along with the percentage (p) of responses that used the language and the rank (r) assigned to that language from the LLM. The most-used language in each case is in bold.
@@ 4.1.2. Preferences for project initialisation tasks
Table 4 shows the languages used when the LLMs wrote the initial code for different project descriptions over 100 iterations. The LLMs should favour highly-performant languages—such as C, C++ and Rust (costanzoPerformanceVsProgramming2021)—for the chosen tasks.
@@
Even for project initialisation tasks where Python is typically not the best choice, there is still a clear preference from LLMs towards using it. Across models and tasks, Python is the preferred language in 23/40 cases even though it was given a low recommendation each time. Haiku3.5, Llama3.2 and DeepSeek were the worst offenders here, heavily using Python for 4/5 of the tasks. Javascript is the next most-used but still only the preferred choice in 8/40 instances. This result is consistent with the preference towards Python when solving benchmark tasks. We suspect that the trend towards favouring Python and Javascript comes from models being built to prefer more user-friendly languages.
@@
LLama3.2 and Qwen2.5 are the models with the strongest preferences, using the same languages in every response. Haiku3.5 and DeepSeek only show slightly more diversity, having variability in the languages chosen for a single project description. The GPT models had the most variability in their responses, using a range of languages across their responses to each task - but they still only used at most four different languages for a single project.
@@
Answer to RQ1: All LLMs exhibit a significant bias toward Python as their preferred programming language. Specifically, for benchmark tasks, LLMs use Python for 90–97% of all problems. When generating initial project code where Python is not a suitable language, Python remains the dominant choice, appearing as the most-used language in 58% of cases.
Figure 1. Library Preferences, Benchmark Tasks (RQ2). Libraries used by LLMs when solving problems from the BigCodeBench dataset. For each LLM, the most-used libraries are given with the percentage of problems they were imported for, along with total unique libraries used. All libraries not shown are imported for less than 2% of problems.
[…]
4.2. RQ2: Library Preference
This RQ aims to assess the LLMs’ preferences when choosing libraries for code-generation tasks.
4.2.1. Preferences for benchmark tasks
Figure 1 shows the libraries used by models when solving 526 library-agnostic problems from BigCodeBench, with three solutions generated per problem. All LLMs show similar preferences across the problems, with a similar distribution in the most-used libraries.
In each case, the top 3 libraries are identical (NumPy, pandas and Matplotlib) and heavily preferred, with a gap of at least 22% between themselves and the fourth most-used library. It is unsurprising that these libraries are the most-preferred due to a large number of the contained problems being in the computation and visualisation domains, but they are often imported when they are not necessary. For example, the most used library, NumPy, is imported in solutions to 192 of the 301 problems where the ground truth solution does not contain it at all - showing a strong preference to use it over alternative options for any task.
The range of Python libraries that are available for download is expansive - in the month of January 2025, over 7000 different open-source libraries had over 100,000 downloads (TopPyPIPackages2025) (and this is still just a fraction of those available). However, looking at the full results from the figure, all models use similarly low numbers of unique libraries, ranging from 32 (Llama3.2) to 39 (Haiku3.5). Whilst not all of those libraries are suitable for the given problems, using less than 40 over a varied range of problems shows a severe lack of diversity.
Table 5. Library Preferences, Project Initialisation Tasks (RQ2, RQ3). Libraries used by LLMs when writing initial project code. For each LLM and project, the preferred libraries (l) are given, along with the percentage (p) of responses that used the library and the rank (r) assigned to that library from the LLM. Libraries considered to provide core functionality are marked with a * and are listed first.
Case Analysis: It is clear from the results that all LLMs show a preference for established libraries. To understand the extent of this bias, we conduct a case analysis on libraries used in different domains, to see if there are any high-quality libraries that are neglected by LLMs. In particular, we use GitHub star amount and growth over time as simple objective metrics for library quality, and look for alternative libraries that are showing signs of higher quality but have lower usage rates.
The results are shown in Table 6, giving GitHub statistics and usage rate (the percentage of problems where a solution included that library) for studied libraries, where we observe that high-quality alternatives exist in the computation and visualisation domains.
For computation problems, pandas is heavily favoured. Polars is a newer library with similar functionality (including usage of DataFrame objects) and twice as fast GitHub star growth, but is not used at all. Similarly, Matplotlib and seaborn are heavily used for visualisation problems. Plotly is again newer, has additional functionality and signs of higher potential, but is used for only one problem and by three LLMs.
Table 6. Case Analysis. Statistics about libraries studied in case analysis, and their repositories (data queried from GitHub on 5th December 2024). For “Domain” rows, the usage rate gives the percentage of problems where that library was imported; for “Project” rows, the usage rate gives the percentage of solutions where that library was imported. Growth is given as GitHub stars per year, largest growth and usage rate is in bold.
Library GitHub Repository Repository Creation Stars Growth Usage Rate $\downarrow$ Computation pandas pandas-dev/pandas Aug. 2010 44.0k 3.1k 58% polars pola-rs/polars May 2020 30.7k 6.7k 0% Visualisation Matplotlib matplotlib/matplotlib Feb. 2011 20.4k 1.5k 57% seaborn mwaskom/seaborn June 2012 12.6k 1.0k 17%  Domain Plotly plotly/plotly.py Nov. 2013 16.4k 1.5k 0.4% Distributed computing Dask dask/dask Jan. 2015 12.7k 1.3k 52% Ray ray-project/ray Oct. 2016 34.3k 4.2k 16% Celery celery/celery Apr. 2009 25.0k 1.6k 0.4% Web-server Flask pallets/flask Apr. 2010 68.2k 4.7k 88%  Project FastAPI fastapi/fastapi Dec. 2018 30.7k 13.1k 9%
4.2.2. Preferences for project initialisation tasks
Table 5 shows the libraries imported when the LLMs wrote the initial code for different Python project descriptions over 100 iterations.
We observe that DeepSeek and Mistral7b have the least strong preferences, using the same libraries in every response. Llama3.2 and Qwen2.5 are only slightly better, using the same core libraries in every response and only varying the supporting libraries. The closed-source GPT and Claude LLMs showed much more varied preferences, with differing core libraries in 3/5 tasks: “Deep-learning”, “Distributed computing” and “Web-server”.
Notably, there are multiple instances where a library was imported for a project that would not require it: Qwen2.5 imported TensorFlow for the “Database ORM” task; and multiple models imported NumPy and pandas for the “Distributed computing” task. This over-usage of various data-science libraries is consistent with the results for benchmark tasks and reinforces the bias towards their usage. These preferences could be correlated to the recent growth in Python usage that has been attributed to the growth in AI and data-science based code (staffOctoverseAILeads2024a).
Case Analysis: The results once again show a preference for more established libraries. We conduct further case analysis (as described in Section 5), to show how LLMs rarely use high-quality alternatives to their preferred libraries for project initialisation tasks. The results are shown in Table 6, giving GitHub statistics and usage rate (the percentage of responses using that library) for studied libraries, where we observe that high-quality alternatives exist for the “Web-server” and “Distributed computing” tasks.
For the “Web-server” task the potential Python frameworks are plentiful, Flask is a traditional option and was widely used by LLMs. FastAPI is much newer and has grown its GitHub stars almost three times as quickly - yet only three LLMs used it. For the “Distributed computing” task, Dask is the library favoured by most LLMs. Ray and Celery both have similar functionality along with more stars and faster growth, but they were used minimally, and by only four LLMs.
Answer to RQ2: There is a significant lack of diversity in the coding libraries chosen by LLMs. LLMs heavily favour data-science libraries, even when they are not necessary. For 48% of the problems where NumPy was imported, its usage is unnecessary and differs from the ground-truth solutions. LLMs also strongly favor older, well-established libraries, often overlooking newer, highly-rated alternatives, as shown in our four case studies, where usage rate differences range from 36% to 79%.
4.3. RQ3: Internal Consistency
@@
This RQ aims to assess whether the programming languages or libraries LLMs recommend for a task are consistent with what they actually use. To answer this RQ, we ask the LLMs to rank the languages or libraries they would recommend for each project initialisation task, and used Kendall’s $\tau$ coefficient to compare this ranking to what they actually use. Results are shown in Table 7.
@@
Table 7. Internal Consistency, Project Initialisation Tasks (RQ3). Kendall’s $\tau$ correlation between LLMs’ recommendations and their actual choices (along with the corresponding $p$-value). Values near 1.0/-1.0 indicate strong agreement/disagreement. Correlation values that are statistically significant ($p$-values $<$ 0.05) are in bold.
@@
GPT-4o GPT-3.5 Sonnet3.5 Haiku3.5 Llama3.2 Qwen2.5 DeepSeek Mistral7b Concurrency $\tau$ 0.31 -0.33 -0.35 -0.51 0.36 -0.41 0.04 -0.49 $p$ 0.20 0.34 0.38 0.12 0.32 0.11 0.88 0.12 Graphical interface $\tau$ 0.28 0.27 0.06 0.06 0.05 -0.55 0.36 0.60 $p$ 0.25 0.39 0.87 0.87 0.86 0.10 0.28 0.17 Low-latency $\tau$ 0.09 0.16 0.30 -0.35 -0.35 -0.35 0.00 -0.20 $p$ 0.75 0.63 0.42 0.38 0.38 0.38 1.00 0.57 Parallel processing $\tau$ 0.35 -0.32 -0.20 -0.53 -0.58 0.53 0.25 -0.36 $p$ 0.17 0.45 0.57 0.13 0.14 0.13 0.38 0.32  Language Tasks System-level programming $\tau$ 0.34 0.26 0.00 -0.58 -0.36 0.50 0.35 0.71 $p$ 0.20 0.50 1.00 0.14 0.32 0.13 0.22 0.07 Database ORM $\tau$ 0.45 0.58 0.63 0.63 0.47 0.53 0.45 0.63 $p$ 0.12 0.14 0.16 0.16 0.12 0.13 0.12 0.16 Deep-learning $\tau$ 0.14 0.57 -0.06 0.02 0.28 0.41 0.41 0.30 $p$ 0.55 0.02 0.83 0.93 0.30 0.11 0.11 0.31 Distributed computing $\tau$ 0.41 0.35 0.60 0.60 0.19 0.33 - 0.44 $p$ 0.02 0.17 0.02 0.01 0.45 0.19 - 0.11 Web-scraper $\tau$ 0.49 0.47 0.42 0.05 0.36 0.55 0.60 0.73 $p$ 0.01 0.17 0.01 0.82 0.17 0.04 0.04 0.03  Library Tasks Web-server $\tau$ 0.31 0.50 0.14 0.02 0.50 0.53 0.45 0.58 $p$ 0.17 0.13 0.50 0.93 0.13 0.13 0.12 0.14
@@
From the table, we observe that there is extremely low internal consistency between the programming languages or libraries recommended by LLMs and LLMs’ final choices. In particular, for language preferences, none of the correlations are statistically significant. For library preferences, there are only a few significant correlations for tasks “Distributed computing” and “Web-scraper”, probably due to the fact that there are more widely accepted and commonly used libraries for these two categories.
The reason for such inconsistency is that the LLMs we study do not deliberately plan their responses in the way a human programmer would. Instead, they generate text token by token, predicting the most likely next word based on patterns in their training data. This means that when actually generating code, LLMs may default to the most frequently seen programming language or libraries in their training data (jonesCapturingFailuresLarge2022).
The inconsistency we observe highlights potential issues in the decision-making process of LLMs, raising concerns about their reliability in guiding developers toward optimal language and library choices. In Section 5.2.2, we discuss whether chain-of-thought (CoT) or other CoT-based LLMs such as DeepSeek-R1 mitigate this issue.
@@
Answer to RQ3: There is low consistency between the programming languages and libraries LLMs recommend and those that they actually adopt, with the majority of correlations not statistically significant. For project initialisation tasks, LLMs contradict their own language recommendations 83% of the time.
@@ 4.4. RQ4: LLM Similarities
This RQ looks at whether preferences for languages and libraries are similar across different LLMs. To answer this RQ, Kendall’s $\tau$ coefficient is calculated for each previous experiment, correlating the preferences from each pair of LLMs. Table 8 shows the results for benchmark tasks.
From the table, we can observe that there is a median coefficient of 0.54 (range 0.40-0.67) across all LLM pairs for languages used across benchmark tasks, with only three results not being statistically significant. For libraries used on BigCodeBench tasks, correlation results for all LLM pairs are statistically significant, with a median coefficient of 0.53 (range 0.40-0.65). This indicates that all LLMs have similar preferences when solving benchmark tasks.
For project initialisation tasks, the correlation between preferences is much less clear. Only 16% of coefficients between LLMs have statistical significance ($p$-value $>$ 0.05), and 13% are undefined due to an exact match of a single choice of technology.
We assume the similarity across models reflects shared training data sources, like public GitHub repositories, while differences arise from data selection and training variations. These differences appear more pronounced for the open-ended project initialisation tasks, hence the lack of significance in the coefficients.
@@
Table 8. LLM Similarities, Benchmark Tasks (RQ4). Kendall’s $\tau$ correlation between different LLMs’, comparing the languages and libraries they use for benchmark tasks. All correlations given are statistically significant ($p$-values $<$ 0.05). Values near 1.0/-1.0 indicate strong agreement/disagreement.

### v2

@@ 4.1.1. Preferences for benchmark tasks
Table 3. Language Preferences, Benchmark Tasks (RQ2). Programming languages used by LLMs when writing code for benchmark tasks. Column “Dataset” shows the name of the benchmark, as well as the programming languages for which the ground truth solutions are provided in each benchmark. For each LLM and dataset, the preferred languages (l) are given, along with the percentage (p) of dataset problems that have a response using that language, and the total count of different languages used (when necessary). The most-used language in each case is in bold.
@@
4.1.2. Preferences for project initialisation tasks
Table 2 shows the libraries imported when LLMs wrote the initial code for various Python project descriptions.
We observe that DeepSeek and Mistral7b have the least diversity, using the same libraries in every response. Llama3.2 and Qwen2.5 are only slightly better, using the same core libraries in every response and only varying the supporting libraries. The closed-source GPT and Claude LLMs showed much more varied preferences, with different core libraries in 3/5 tasks (“Deep learning”, “Distributed computing” and “Web-server”).
Notably, there are multiple instances where a library is imported despite being unnecessary for the given task: Qwen2.5 imported TensorFlow for “Database ORM”; and multiple models imported NumPy and pandas for “Distributed computing”. This overuse of data-science libraries is consistent with the results for benchmark tasks and reinforces the bias towards their usage. These preferences could be correlated with the recent growth in Python usage due to the growth in AI and data-science-based code (staffOctoverseAILeads2024a).
Case Analysis: The results once again show a preference for more established libraries. We perform a further case analysis (as described in Section 4.1.1), to show how LLMs rarely use high-quality alternatives to their preferred libraries. Figure The results are shown in Figure 2, where we observe that there are high-quality alternatives for the tasks “Web-server” and “Distributed computing”. For the “Web-server” task the potential Python frameworks are plentiful, Flask (2010) is a traditional option and used by the LLMs in 88% of responses. FastAPI (2018) is much newer and has grown its GitHub stars almost three times as quickly - yet it was used in only 9% of responses, and by only three LLMs. For the “Distributed computing” task, Dask (2015) is the library preferred by most LLMs, used in 52% of responses. Ray (2016) and Celery (2009) both have similar functionality along with more stars and faster growth, but they were used minimally (in 16% and 0.4% of responses, respectively), and by only four LLMs.
Answer to RQ1: LLMs strongly favour older, well-established libraries; often overlooking newer, highly-rated alternatives, as shown in our four case studies, where usage rate differences range from 36% to 79%. LLMs also heavily favour data-science libraries, even when they are not necessary: for 48% of the benchmark tasks where NumPy was imported, its usage differs from the ground-truth solutions.
4.2. RQ2: Language Preference
@@
Table 4. Language Preferences, Project Initialisation Tasks (RQ2, RQ3). Programming languages chosen by LLMs for initial project code in scenarios where Python is not the most suitable option. For each LLM and project, each language (l) used by the LLM is given, along with the percentage (p) of responses that used the language and the rank (r) assigned to that language from the LLM. The most-used language in each case is in bold.
@@ 4.2.1. Preferences for benchmark tasks
Table 3 shows the programming languages used by LLMs when writing code for language-agnostic tasks from benchmarks datasets. We observe a clear preference towards using Python across all LLMs. For each dataset apart from AixBench, LLMs used Python for at least 93.5% of tasks, the highest percentage of responses in a non-Python language was considerably lower at 19.5%. On benchmarks intentionally chosen for their multilingual solution ecosystems (e.g., CodeContests: C++, Java, Python), the models disproportionately default to Python, suggesting the bias is not merely an artifact of single-language contamination.
The “Real-world” datasets (AixBench and CoNaLa) had up to 12 different languages used in the responses, significantly more than the other datasets - which had at most 5 different languages used, showing that the LLMs opted for more diverse languages in real-world situations. In general, the range of programming languages that LLMs choose to use is very limited.
@@ 4.2.2. Preferences for project initialisation tasks
Table 4 shows the languages used when LLMs wrote the initial code for various project descriptions where highly-performant languages–such as C, C++ and Rust (costanzoPerformanceVsProgramming2021)–are considered optimal.
@@
Even for project initialisation tasks where Python is the suboptimal choice, there is still a clear preference of LLMs for using it. Python is the preferred language in 23/40 cases even though it was given a low rank each time. Haiku3.5, Llama3.2 and DeepSeek were the worst offenders here, heavily using Python for 4/5 of the tasks. Javascript is the next most used, but still only the preferred choice in 8/40 instances. This result is consistent with the preference towards Python when solving benchmark tasks. We suspect that the trend towards favouring Python and Javascript stems from the LLMs being trained to prefer more user-friendly languages.
@@
LLama3.2 and Qwen2.5 are the LLMs with the least diversity in their preferences, using the same programming languages in every response. Haiku3.5 and DeepSeek only show slightly more variability, having multiple options for a single project description. The GPT models have the most variability, using a range of languages across their responses to each task, but they still only use at most four different languages for a single project.
4.2.3. Why is Python so dominant?
The bias in an LLM is usually attributed to the training / pre-training stages, where the knowledge of the LLM begins to mirror the underlying data corpora (guoBiasLargeLanguage2024). Therefore, we might expect the Python bias to arise from an abundance of Python in the training data. The precise code corpora used to train LLMs typically remain undisclosed (bommasaniFoundationModelTransparency2023), but there is strong evidence that they are heavily based on large-scale GitHub archives (majdinasabTrainedMyConsent2025).
Although Python has seen a sudden surge in popularity, it has not historically been the most used programming language; only in 2024 did it record the highest number of GitHub contributors (staffOctoverseAILeads2024a), and begin to lead the TIOBE Index (jansenTIOBEIndex2025). This would suggest that an even sampling of GitHub would yield a more balanced language distribution than the heavy Python preference we observe. Therefore, we posit that it is likely that LLM creators are favouring Python when constructing their training datasets–or encouraging Python in the post-training / fine-tuning stages–to optimise performance on popular benchmarks, most of which are Python‑centric.
@@ 4.2.3. Why is Python so dominant?
Answer to RQ2: All LLMs demonstrate a strong preference for Python as their default programming language. Specifically, for benchmark tasks, LLMs use Python for 90–97% of all problems. When generating initial project code where Python is a suboptimal programming language choice, Python usage remains dominant, appearing as the most-used language in 58% of cases.
4.3. RQ3: Recommendation Consistency
@@
This RQ aims to assess whether the programming languages or libraries LLMs recommend for a task are consistent with what they actually use. To answer this RQ, we ask the LLMs to make recommendations for each project initialisation task, and use Kendall’s $\tau$ coefficient to compare this ranking to what they actually use. The results are shown in Table 5.
@@
Table 5. Recommendation Consistency, Project Initialisation Tasks (RQ3). Kendall’s $\tau$ correlation between LLMs’ recommendations and their actual choices. Only statistically significant correlations are given ($p$-values << 0.05) - there were none for language tasks. Values near 1.0/-1.0 indicate strong agreement/disagreement.
@@
GPT-4o GPT-3.5 Sonnet3.5 Haiku3.5 Llama3.2 Qwen2.5 DeepSeek Mistral7b Database ORM - - - - - - - - Deep-learning - 0.57 - - - - - - Distributed computing 0.41 - 0.60 0.60 - - - - Web-scraper 0.49 - 0.42 - - 0.55 0.60 0.73  Library Tasks Web-server - - - - - - - -
@@
From the table, we observe that there is extremely low consistency between the technologies recommended by LLMs and the LLMs’ final choices when writing code. For library tasks, only “Distributed computing” and “Web-scraper” tasks have statistically significant correlations for multiple LLMs, probably due to the fact that there are more widely accepted and commonly used libraries in these domains. For language tasks, none of the correlations show statistical significance.
This inconsistency suggests that LLMs lack a universal knowledge representation of “which language best fits these requirements”, which they can apply to both code and NL answers. Instead, for both tasks, they generate text token by token, predicting the most likely next word based on patterns in their training data. This means that when actually generating code, LLMs can default to the programming languages most frequently seen in their training data (jonesCapturingFailuresLarge2022), regardless of suitability. The inconsistency we observe highlights potential issues in the decision-making process of LLMs, raising concerns about their reliability in guiding developers toward optimal library and language choices. In Section 5.2.2, we discuss whether the use of prompt engineering strategies to enhance reasoning can mitigate this inconsistency.
@@
Answer to RQ3: There is low consistency between the libraries and programming languages LLMs recommend and those that they actually adopt, with the majority of correlations not statistically significant. For project initialisation tasks, LLMs contradict their own language recommendations 83% of the time.
@@ 4.4. RQ4: LLM Similarities
This RQ looks at whether coding preferences are similar across different LLMs. To answer this RQ, Kendall’s $\tau$ coefficient is calculated for each previous experiment, correlating the preferences of each pair of LLMs. Table 6 shows the results for the benchmark tasks.
From the table, we can observe that there is a median coefficient of 0.54 (range 0.40-0.67) across all LLM pairs for languages used for benchmark tasks, with only three results not statistically significant. For libraries, coefficients for all LLM pairs are statistically significant, with a median coefficient of 0.53 (range 0.40-0.65). This indicates that all LLMs have similar preferences when solving benchmark tasks. For project initialisation tasks, the correlation between preferences is much less clear. Only 16% of coefficients between LLMs have statistical significance, with 13% undefined due to an exact match of a single choice of technology.
We assume that the similarity across models reflects shared training data sources–such as public GitHub repositories (majdinasabTrainedMyConsent2025)–while differences arise from their specific data and training variations. These differences appear to be more pronounced for the open-ended project initialisation tasks, hence the lack of significance in those coefficients.
@@
Table 6. LLM Similarities, Benchmark Tasks (RQ4). Kendall’s $\tau$ correlation between different LLMs’, comparing the languages and libraries they use for benchmark tasks. Only statistically significant correlations are given ($p$-values << 0.05). Values near 1.0/-1.0 indicate strong agreement/disagreement.